# M.J. Lopez-Herrero ✉ 11 $0$-0003-2835-2940    D. Taipe 11 $0$-0002-2207-4129

### v1

@@ 1 Introduction
In this paper, we consider an infectious disease taking place in a finite and isolated population. We assume that the disease is transmitted by direct contact with an infectious individual. After recovery, individuals show temporary immunity. In addition, we assume that the disease is a vaccine-preventable infection, but vaccinated individuals are not lifelong fully protected, either because of vaccine failures or because of vanning immunity. Given the latter fact, re-vaccination of susceptible individuals is planned as the epidemic progresses.
@@ 2.1 Long term behavior
$p_{s}=\binom{N}{s}(\frac{1}{1+\rho/\theta})^{s}(\frac{\rho/\theta}{1+\rho/\theta})^{N-s},\text{ for }0\leq s\leq N.$
@@
$lim_{t\xrightarrow{}\infty}P\{I(t)=i,S(t)=s,V(t)=v\}=\delta_{i,0}\delta_{v,N-s}\binom{N}{s}(\frac{\theta}{\theta+\rho})^{s}(\frac{\rho}{\theta+\rho})^{v},$
@@ 3 Incidence of the disease during an outbreak
Note that the right-hand side term of equation (LABEL:momat) agrees with the definition of $D^{k}_{i}$ in the statement of Theorem 3.2.
Finally, a standard forward elimination backward substitution procedure gives the expressions (23) and (24).
@@ 4 Numerical Results
Using the explicit value of $E[V]$ along with the vanning rate of the vaccine, for a given contact disease and an appropriate vaccination coverage $v_{c}=Nf$, our criterion provides the booster rate $\rho$ such that $E[V]\geq v_{c}$. That is,
@@
For example, for the Ebola virus, the transmission $R_{0}$ is around 1.2. Vaccines have only recently been introduced and there is not yet enough information on the waning effects, but their efficacy varies from 75% to 100%. So, in an isolated population of 100 people, we need to keep at least 18 vaccinated if the vaccine is 95% effective, or at least 21 if the vaccine is 80% effective. Correspondingly, to maintain the herd protection level in the long term, the ratio between booster doses and waning time is 4.55 (i.e.; the mean time between booster doses is at most 4.55 the expected duration of vaccine protection) or 3.76 (if the vaccine is 80% effective). For a disease with basic reproduction number $R_{0}$ close to 4.2, such as diphtheria, if we administer a 95% effective vaccine, we observe that $v_{c}=81$ and $\rho/\theta=0.23$. Therefore, if we fix the number of vaccine protected individuals to be at least 81 in the long run, the mean time between booster doses should be approximately the fourth part of the expected duration of vaccine protection.
@@ 5 Conclusions
In this paper we consider a stochastic SIVRS model with imperfect vaccine and loss of protection. The mathematical model used to describe how the epidemic evolves is a compartmental stochastic model involving a three-dimensional CTMC. The finite size and isolation of the population guarantees that the pathogen will disappear as time goes on. And the stationary distribution gives chance to situations where only vaccinated and susceptible individuals exist in the population. The form of this distribution is binomial of N trials and probability of success depending on the quotient between the waning effect rate and the vaccination rate.

### v2

@@ 1 Introduction
In this paper, we consider an infectious disease taking place in a finite and isolated population. We assume that the disease is transmitted by direct contact with an infectious individual. After recovery, individuals show temporary immunity. In addition, we assume that the disease is a vaccine-preventable infection, but vaccinated individuals are not lifelong fully protected, either because of vaccine failures or because of waning immunity. Given the latter fact, re-vaccination of susceptible individuals is planned as the epidemic progresses.
@@ 2.1 Long term behavior
$p_{s}=\binom{N}{s}\left(\frac{1}{1+\rho/\theta}\right)^{s}\left(\frac{\rho/\theta}{1+\rho/\theta}\right)^{N-s},\text{ for }0\leq s\leq N.$
@@
$lim_{t\xrightarrow{}\infty}P\{I(t)=i,S(t)=s,V(t)=v\}=\delta_{i,0}\delta_{v,N-s}\binom{N}{s}\left(\frac{\theta}{\theta+\rho}\right)^{s}\left(\frac{\rho}{\theta+\rho}\right)^{v},$
@@ 3 Incidence of the disease during an outbreak
Note that the right-hand side term of equation (LABEL:momat) agrees with the definition of ${\bf D}^{k}_{i}$ in the statement of Theorem 3.2.
Finally, a standard forward elimination-backward substitution procedure gives the expressions (23) and (24).
@@ 4 Numerical Results
Using the explicit value of $E[V]$ along with the waning rate of the vaccine, for a given contact disease and an appropriate vaccination coverage $v_{c}=Nf$, our criterion provides the booster rate $\rho$ such that $E[V]\geq v_{c}$. That is,
@@
For example, for the Ebola virus, the transmission $R_{0}$ is around 1.2. Vaccines have only recently been introduced and there is not yet enough information on the waning effects, but their efficacy varies from 75% to 100%. So, in an isolated population of 100 people, we need to keep at least 18 vaccinated if the vaccine is 95% effective, or at least 21 if the vaccine is 80% effective. Correspondingly, to maintain the herd protection level in the long term, the ratio between booster doses and waning time is 4.55 (i.e.; the mean time between booster doses ($1/\rho$) is at most 4.55 the expected duration of vaccine protection ($1/\theta$)) or 3.76 (if the vaccine is 80% effective). For a disease with basic reproduction number $R_{0}$ close to 4.2, such as diphtheria, if we administer a 95% effective vaccine, we observe that $v_{c}=81$ and $\theta/\rho=0.23$. Therefore, if we fix the number of vaccine protected individuals to be at least 81 in the long run, the mean time between booster doses should be approximately the fourth part of the expected duration of vaccine protection.
@@ 5 Conclusions
In this paper we consider a stochastic SVIRS model with imperfect vaccine and loss of protection. The mathematical model used to describe how the epidemic evolves is a compartmental stochastic model involving a three-dimensional CTMC. The finite size and isolation of the population guarantees that the pathogen will disappear as time goes on. And the stationary distribution gives chance to situations where only vaccinated and susceptible individuals exist in the population. The form of this distribution is binomial of N trials and probability of success depending on the quotient between the waning effect rate and the vaccination rate.
@@ 6 Acknowledgments
7 Appendix: Supplemental material
In this Appendix, we characterize the infinitesimal generator ${\bf Q}$ of the three-dimensional CTMC ${\cal X}$ with state space ${\cal S}$
$\mathcal{S}=\{(i,s,v):0\leq i,s,v\leq 0,\quad i+s+v\leq N\}.$
Note that $\mathcal{S}$ containing $\binom{N+3}{3}$ states, that can be organized into levels according to the number of infected individuals, as follows:
${\cal S}=\cup_{i=0}^{N}S(i)$
where, for $0\leq i\leq N$, each set $S(i)$ is partitioned in terms of sub-levels as $S(i)=\cup_{s=0}^{N-i}l(i,s)$ with $l(i,s)=\{(i,s,v):0\leq v\leq N-i-s\}$. In particular, the order of sub-levels inside $S(i)$ is
$l(i,0)\prec l(i,1)\prec\ldots\prec l(i,N-i),$
while we order the states in $l(i,s)$ according to the number of vaccine protected individuals. That is
$(i,s,0)\prec(i,s,1)\prec\dots\prec(i,s,N-i-s).$
Thus, the number of states in each level $S(i)$ is given by #$\#S(i)=\binom{N-i+2}{2}$ and each sub-level $l(i,s)$ contains #$\#l(i,s)=N-i-s+1$, for $0\leq i\leq N$ and $0\leq s\leq N-i$.
The above partition of the state space ${\cal S}$ yields a block-tridiagonal structure for the infinitesimal generator of ${\cal X}$ that looks as follows:
${\bf Q}$ $=$ $\left(\begin{array}[]{ccccc}{\bf A}_{0,0}&{\bf A}_{0,1}&&&\\ {\bf A}_{1,0}&{\bf A}_{1,1}&{\bf A}_{1,2}&&\\ &\ddots&\ddots&\ddots&\\ &&{\bf A}_{N-1,N-2}&{\bf A}_{N-1,N-1}&{\bf A}_{N-1,N}\\ &&&{\bf A}_{N,N-1}&{\bf A}_{N,N}\end{array}\right),$
where the entries of the sub-matrices $A_{i,i*}$ are the transition rates from states in $S(i)$ to states in $S(i*)$. Specifically, the non-null elements are described as follows:
  • For $1\leq i\leq N$, sub-matrix $\mathbf{A}_{i,i-1}$ is associated with jumps of process ${\cal X}$ from states in level $S(i)$ to states in level $S(i-1)$. In more detail, it has the following form
    ${\bf A}_{i,i-1}$ $=$ $\left(\begin{array}[]{ccccc}{\bf B}_{i,i-1}(0,0)&&&&\\ &{\bf B}_{i,i-1}(1,1)&&&\\ &&\ddots&&\\ &&&{\bf B}_{i,i-1}(N-i,N-i)&0\end{array}\right).$
    Matrices $\mathbf{B}_{i,i-1}(s,s)$ record transition rates from sub-level $l(i,s)$ to sub-level $l(i-1,s)$. We express the particular case of $i=N$ as
    ${\bf A}_{N,N-1}$ $=$ $\left(\begin{array}[]{cc}{\bf B}_{N,N-1}(0,0),&0\end{array}\right).$
  • For $0\leq i\leq N$, sub-matrix $\mathbf{A}_{i,i}$ is related to transitions from level $S(i)$ to level $S(i)$, and is given by
    $\mathbf{A}_{i,i}=\left(\begin{array}[]{ccccc}{\bf B}_{i,i}(0,0)&{\bf B}_{i,i}(0,1)&&&\\ {\bf B}_{i,i}(1,0)&{\bf B}_{i,i}(1,1)&{\bf B}_{i,i}(1,2)&&\\ &\ddots&\ddots&\ddots&\\ &&{\bf B}_{i,i}(N-i-1,N-i-2)&{\bf B}_{i,i}(N-i-1,N-i-1)&{\bf B}_{i,i}(N-i-1,N-i)\\ &&&{\bf B}_{i,i}(N-i,N-i-1)&{\bf B}_{i,i}(N-i,N-i)\end{array}\right).$
    Matrices ${\bf B}_{i,i}(s,s^{\prime})$ record transition rates from sub-level $l(i,s)$ to sub-level $l(i,s^{\prime})$, for $s-1\leq s^{\prime}\leq s+1$. It should be pointed out the particular case $i=N$ which leads to
    ${\bf A}_{N,N}=\left(-q(N,0,0)\right),$
    where $q(N,0,0)=\gamma N$.
  • For $10\leq i\leq N-$ 1, sub-matrix $\mathbf{A}_{i,i+1}$, corresponds to transitions from states in level $S(i)$ to states in level $S(i+1)$, and is given by
    $\mathbf{A}_{i,i+1}=\left(\begin{array}[]{ccccc}{\bf B}_{i,i+1}(0,0)&&&&\\ {\bf B}_{i,i+1}(1,0)&{\bf B}_{i,i+1}(1,1)&&&\\ &{\bf B}_{i,i+1}(2,1)&{\bf B}_{i,i+1}(2,2)&&\\ &&\ddots&\ddots&\\ &&&{\bf B}_{i,i+1}(N-i-1,N-i-2)&{\bf B}_{i,i+1}(N-i-1,N-i-1)\\ &&&&{\bf B}_{i,i+1}(N-i,N-i-1)\end{array}\right),$
    where matrices ${\bf B}_{i,i+1}(s,s^{\prime})$ are related to transition rates from sub-level $l(i,s^{\prime})$ to sub-level $l(i+1,s^{\prime})$, for $s^{\prime}=s-1,s$.
The above matrices ${\bf B}_{i,i^{\prime}}(s,s^{\prime})$, for $i-1\leq i^{\prime}\leq i+1$, and $s-1\leq s^{\prime}\leq s+1$, are described below.
  • For $1\leq i\leq N$ and $0\leq s\leq N-i$, matrix $\mathbf{B}_{i,i-1}(s,s)$ contains transitions rates from states in level $l(i,s)$ to states in level $l(i-1,s)$, and is given by
    $\mathbf{B}_{i,i-1}(s,s)=\left(\gamma i{\bf I}_{N-i-s+1}\quad{\bf 0}_{N-i-s+1}\right),$
    where ${\bf I}_{a}$ is the identity matrix of order $a$.
  • For $0\leq i\leq N$ and $1\leq s\leq N-i$, matrix ${\bf B}_{i,i}(s,s-1)$ contains transitions rates from states in level $l(i,s)$ to states in level $l(i,s-1)$, and is given by
    $\mathbf{B}_{i,i}(s,s-1)=\left({\bf 0}_{N-i-s+1}\quad\rho s{\bf I}_{N-i-s+1}\right).$
  • For $0\leq i\leq N$ and $0\leq s\leq N-i$, matrix $\mathbf{B}_{i,i}(s,s)$ contains transitions rates from states in level $l(i,s)$ to states in level $l(i,s)$, and is given by
    $\mathbf{B}_{i,i}(s,s)=diag\left(-q_{(i,s,v)}:0\leq v\leq N-i-s\;\right),$
    with $q(i,s,v)=\beta(i,s)+\eta(i,v)+\gamma(i)+\rho s+\theta v+\epsilon(N-i-s-v)$, where $diag(a_{1},a_{2},...,a_{m})$ is a diagonal matrix with non-null entries $a_{1},a_{2},...,a_{m}$.
  • For $0\leq i\leq N$ and $10\leq s\leq N-i-$ 1, matrix $\mathbf{B}_{i,i}(s,s+1)$ contains transitions rates from states in level $l(i,s)$ to states in level $l(i,s+1)$, and is given by
    $\mathbf{B}_{i,i}(s,s+1)=\left(\begin{array}[]{cccccc}\epsilon(N-i-s)&&&&&\\ \theta&\epsilon(N-i-s-1)&&&&\\ &2\theta&\epsilon(N-i-s-2)&&&\\ &&\ddots&\ddots&&\\ &&&(N-i-s-2)\theta&2\epsilon&\\ &&&&(N-i-s-1)\theta&\epsilon\\ &&&&&(N-i-s)\theta\\ \end{array}\right).$
  • For $10\leq i\leq N-$ 1 and $11\leq s\leq N-i-$ 1, matrix $\mathbf{B}_{i,i+1}(s,s-1)$ contains transitions rates from states in level $l(i,s)$ to states in level $l(i+1,s-1)$, and is given by
    $\mathbf{B}_{i,i+1}(s,s-1)=\left(\beta is\;{\bf I}_{N-i-s+1}\right).$
  • For $10\leq i\leq N-$ 1 and $10\leq s\leq N-i-$ 1, matrix $\mathbf{B}_{i,i+1}(s,s)$ contains transitions rates from states in level $l(i,s)$ to states in level $l(i+1,s)$, and is given by
    $\mathbf{B}_{i,i+1}(s,s)$ $=$ $\left(\begin{array}[]{c}{\bf 0}^{\prime}_{N-i-s}\\ diag\left(\beta hiv:1\leq v\leq N-i-s\;\right)\end{array}\right).$